2.3 One merchandising company sells two products (A & B). The products sales mixes are and 60% respectively. Their manufacturing variable costs percentages are 40% and 50% respectively. The marketing variable costs are 10% over revenues in each one. The fixed overheads are 88, 000 per the whole period. 40 % Required: 1. Calculate the contribution margin percentage product A B and the Weighted-average 2. Calculate the revenues companys breakeven point and the same per products. Prepare the 3. Calculate the total revenues needed to earn a target operating income of 22,000 . Prepare 4. If the total amount of fixed costs is 88.000, of which 620.000 are traceable and direct to percentage Income Statement Report. the income statement. product A and 44.000 direct traceable relate to product B, while the remaining costs are common fixed costs for the company, calculate the Critical point of product A and the Critical point of product B. Prepare the Segmented Income Statement Report for that situation
Answer #1

1)Contribution margin %= sales -variable costs )/sales= 1-(variable/sales)
A=1-(40%+10%)=50%
B=1-(50%+10%)=40%
Weighted average %=(50%*40%)+(40%*60%)=44%

2) Company breakeven point in revenue=Fixed costs/Weighted average Contribution margin %
=88000/44%
=200000
For product A=200000*40%=80000
Product B=200000*60%=120000
Income statement:
Revenues=200000
Variable costs=44%*200000=88000
Contribution margin=112000
Fixed costs=88000
Net income=24000

3)Target operating income=22000
Contribution margin required=22000+88000=110000
Revenues=110000/(1-44%)
=196428.57
Income statement:
Revenues=196428.57
Variable costs=44%*196428.57=86428.57
Contribution margin=110000
Fixed costs=88000
Net income=22000

4)critical point for Product=20000/50%=40000
Product B=44000/40%=110000
Segment income statement for A:
Revenues=40000
Variable costs=50%*40000=20000
Contribution margin=20000
Fixed costs=20000
Net income=0

Product B:
Revenues=110000
Variable costs=40%*110000=44000
Contribution margin=66000
Fixed costs=44000
Net income=22000